OpenStudy (anonymous):

What is the number of ways to arrange 6 objects from a set of 10 different objects? A. 520 B. 78,400 C. 151,200 D. 210

OpenStudy (swissgirl):

Its 10C6 which is 10!/(6!(10-6)!) = ..

OpenStudy (anonymous):

how do i work the problem out?

OpenStudy (swissgirl):

We use the n Choose R Formula which is \[\frac{n!}{r!(n-r)!}\] In this case n=10 cuz there are 10 options and r=6 cuz thats the amount we are selecting

OpenStudy (anonymous):